OpenStudy (anonymous):

Solve x^3-64=0, There are 3 solutions?

OpenStudy (jackellyn):

Do you know the formula to factor cubes?

OpenStudy (anonymous):

(x-y)(x^2+xy+y^2)...?

OpenStudy (jackellyn):

a^3 + b^3 = (a + b)(a^2 – ab + b^2) a^3 – b^3 = (a – b)(a^2 + ab + b^2)

OpenStudy (jackellyn):

There could be 3 roots but it isnt always the case, to answer your question.

OpenStudy (jackellyn):

You would use the second one so (x-4)(x^2+4x+16) right?

OpenStudy (anonymous):

Yep. I got that. I also did x^3-64=0 x^3=64 and took the cube root of both so... x=4. I guess I don't understand how to get the 3rd solution?

OpenStudy (jackellyn):

x=4 is one and then to find the other 2 you have to use the quadratic formula \[(-b \pm \sqrt{b ^{2}-4ac})\div2\]

OpenStudy (anonymous):

Thanks!